a vector has an x-component of -2.5 meters and a y-component of 4.2 meters. Express the vector in magnitude angle form.
1
Expert's answer
2019-10-17T11:13:19-0400

Let the vector "\bf a" has components

"a_x=-2.5\:\rm m"

"a_y=4.2\:\rm m"

Thus, the magnitude of the vector "\bf a"


"a=\sqrt{a_x^2+a_y^2}=\sqrt{(-2.5)^2+4.2^2}=4.9\:\rm m"

The angle that vector "\bf a" makes with positive x-direction


"\theta=\tan^{-1}\frac{a_y}{a_x}=\tan^{-1}\frac{4.2}{-2.5}=121^{\circ}"

Answer: "{\bf a}=(4.9\:\rm m;121^{\circ})"